The difference between two numbers is 7. The sum of half the larger number and twice the smaller number is 51. What are the two numbers? must set up the equations and show in algebraic steps)
Answer by deborabr(181)

this equations system is representing the problem
x-y=7
x/2+2y=51
.
x-y=7 *(2) multiplying to when you sum delet y
x/2+2y=51
.
2x-2y=14
x/2+2y=51 suming
___________
2x+x/2+0y=65
2x+x/2=65
4x/2+x/2=130/2 make the denominators equal to delete them
4x+x=130
5x=130
x=130/5
x=26
now substituting x to find y
x-y=7
26-y=7
26-7=y
19=y
the numbers are 19 and 26